Dream, Inc., has debt outstanding with a face value of $4 million. The value of the firm if it were entirely financed by equity would be $18.5 million. The company also has 490,000 shares of stock outstanding that sell at a price of $32 per share. The corporate tax rate is 35 percent. What is the decrease in the value of the company due to expected bankruptcy costs?
Answer #1

Calculation of decrease in the value of the company due to expected bankruptcy costs

STEP- I Calculation of levered firm as per MM Proposition-I with taxes

Value of Levered firm = Value of Un levered firm + ( Debt * Tax rate )

=18.5 million + ( 4 million * 35% )

=18.5 + 1.4

=199,00,000

STEP - II Value of Market value of firm

Market value of firm = 40,00,000 + 490,000*32

=196,80,000

So decrease in the value of the company due to expected bankruptcy costs

=199,00,000-196,80,000

=220.000